# DHG-Bench: A Comprehensive Benchmark on Deep Hypergraph Learning

### v1

@@ A.2 Detailed description of algorithms
In our developed DHG-Bench, we integrate 16 state-of-the-art HNN algorithms, including 10 spectral-based methods, 4 spatial-based methods, and 2 tensor-based methods. Specifically, we introduce all the HNN models as follows.
Spectral-based HNN Models
@@
HyperGCN (Yadati et al., 2019): HyperGCN approximates each hyperedge of the hypergraph by a set of pairwise edges connecting the vertices of the hyperedge and treats the learning problem as a graph learning problem on the approximation.
@@
HyperND (Prokopchik et al., 2022): HyperND develops a nonlinear diffusion process on hypergraphs that spreads both features and labels following the hypergraph structure. The novel diffusion can incorporate a broad class of nonlinearities to increase the modeling capability, and the limiting point serves as a node embedding from which we make predictions with a linear model.
@@
SheafHyperGNN (Duta et al., 2023): SheafHyperGNN introduces a cellular sheaf framework for hypergraphs, enabling the modeling of complex dynamics while preserving their higher-order connectivity. Then, it generalize the two commonly used hypergraph Laplacians to incorporate the richer structure sheaves offer and construct two powerful neural networks capable of inferring and processing hypergraph sheaf structure.
HJRL (Yan et al., 2024): HJRL introduces a novel cross expansion method, which transforms both hypervertices and edges of a hypergraph to vertices in a standard graph. Then, a joint learning model is proposed to embed both hypervertices and hyperedges into a shared representation space. In addition, the algorithm employs a hypergraph reconstruction objective to preserve structure information in the model.
@@
DPHGNN (Saxena et al., 2024): DPHGNN is a hybrid spectral-spatial framework designed for effective feature representation in resource-constrained hypergraph settings. It introduces equivariant operator learning to capture lower-order semantics by inducing topology-aware inductive biases. It employs a dual-layered feature update mechanism: a static update layer provides spectral biases and relational features, while a dynamic update layer fuses explicitly aggregated features from the underlying graph topology into the hypergraph message-passing process.
Spatial-based HNN Models
HNHN (Dong et al., 2020): HNHN is a hypergraph convolution network with nonlinear activation functions applied to both hypernodes and hyperedges, combined with a normalization scheme that can flexibly adjust the importance of high cardinality hyperedges and high-degree vertices depending on the dataset.
@@
AllSetTransformer (Chien et al., 2022): AllSet is a novel HNN paradigm that implements HNN layers as compositions of two multiset functions. The AllSetTransformer integrates powerful Set Transformer (Lee et al., 2019) in the design of the specialized HNN, enabling greater modeling flexibility and enhanced expressive power.
ED-HNN (Wang et al., 2023a): ED-HNN is a HNN architecture designed to approximate any continuous, permutation-equivariant hypergraph diffusion operator. The model is efficiently implemented by combining the star expansion (bipartite representation) of hypergraphs with standard message passing neural networks, and supports scalable training via shared weights across layers.
Tensor-based HNN Models
@@
T-HyperGNN (Wang et al., 2024a): T-HyperGNN is a general framework that integrates tensor hypergraph signal processing (t-HGSP) (Pena-Pena et al., 2023) to encode hypergraph structures using tensors. It models node interactions through multiplicative interaction tensors, elevating aggregation from traditional linear operations to higher-order polynomial mappings, thereby enhancing expressive power. To ensure scalability, T-HyperGNN introduces tensor-message-passing by exploiting tensor sparsity, enabling efficient processing of large hypergraphs with computational and memory costs comparable to matrix-based HNNs.
Appendix B Additional Experimental Details
@@ B.1 General experimental settings
| SheafHyperGNN | init_hedge | ’rand’, ’avg’ |
| sheaf_pred_block | ’MLP_var1’, ’MLP_var2’, ’MLP_var3’, ’cp_decomp’ | |
@@
| TFHNN | mlp_hidden_size | 64, 128, 256, 512, 1024 |
| mlp_num_layers | 1, 2, 3, 4 | |
@@
| EDHNN | alpha | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9 |
@@
We strive to follow the original implementations of various HNN methods from their respective papers or source codes and integrate them into a unified training and evaluation framework. All parameters are randomly initialized. We use the cross-entropy loss function (Mao et al., 2023) for all three benchmark classification tasks. Adam optimizer (Kingma, 2014) is adopted with an appropriate learning rate and weight decay to achieve the best performance on the validation split. We repeat all experiments five times and report the average results along with standard deviations.
B.2 Evaluation metrics
We incorporate a variety of metrics to ensure a comprehensive benchmark evaluation across different tasks.
For node classification task, we adopt widely-used Accuracy (Acc) (Feng et al., 2019; Chien et al., 2022; Wang et al., 2025) to measure the predictive capacity of each HNN algorithm. As for fairness evaluation, we utilize two widely used group fairness metrics: demographic parity () (Dwork et al., 2012), and equalized odds () (Hardt et al., 2016). Specifically, is defined as the acceptance rate difference between the two sensitive groups, which can be formulated as:
where denotes predicted node label and is the sensitive attribute. Similarly, quantifies the disparity in predictions by further conditioning on the ground truth , formulated as:
Both metrics are estimated on the test dataset. Lower values of and indicate better fairness.
For hyperedge prediction task, following (Hwang et al., 2022; Ko et al., 2025; Yu et al., 2025), we use AUROC (Area Under the Receiver Operating Characteristic Curve) and AP (Average Precision), where higher values of these metrics indicate stronger hyperedge prediction performance.
For the hypergraph classification task, following (Feng et al., 2024), we report both accuracy and macro-F1 score, where macro-F1 provides a balanced assessment of classification performance across all classes, especially in the presence of class imbalance.
B.3 Hyperparameter setting
We carefully tune hyperparameters to ensure a rigorous and unbiased evaluation of the integrated HNN methods. For algorithms without explicit hyperparameter guidelines in their original papers or source code, we perform a grid search to identify optimal configurations. The search spaces are provided in Table A4. For detailed interpretations, please refer to the corresponding papers. All complete hyperparameter configurations are available in our publicly released GitHub repository.
B.4 Experimental environment
@@
Appendix C Reproducibility
Accessibility. All the datasets, algorithm implementations, experimental settings, and hyper-parameter configurations are publicly available in our open-source library DHG-Bench at https://github.com/Coco-Hut/DHG-Bench.
License. Our package (codes and datasets) is licensed under the MIT License. This license permits users to freely use, copy, modify, merge, publish, distribute, sublicense, and sell copies of the software, provided that the original copyright notice and permission notice are included in all copies or substantial portions of the software. The MIT License is widely accepted for its simplicity and permissive terms, ensuring ease of use and contribution to the codes and datasets. We bear all responsibility in case of violation of rights, etc, and confirmation of the data license.
Code Maintenance. We are committed to continuously updating our code and and actively addressing users’ issues and feedback. Additionally, we warmly welcome community contributions to enhance our library and benchmark algorithms. Nonetheless, we will enforce strict version control measures to ensure reproducibility throughout the maintenance process.
@@ Appendix D Further Discussions
D.1 Related works.
@@
Deep hypergraph learning (DHGL) have emerged as a powerful tool for mining the knowledge of higher-order interactions, which are pervasive in real-world complex systems and applications. Kim et al. (Kim et al., 2024b) recently present the first survey dedicated to HNNs, with an in-depth and step-by-step guide. The survey comprehensively reviews existing HNN architectures, training strategies, and applications, establishing a foundational understanding crucial for advancing the field of DHGL. However, systematic and comprehensive evaluations of HNN algorithms remain scarce. HyFER (Hwang et al., 2021) represents the first well-structured framework for fair HNN algorithm comparison, modularizing the training process into data, model, and task components. Moreover, to address the scalability problem that most existing implementations suffer from, HyFER is built on top of Deep Graph Library (DGL) (Wang et al., 2019), which is a highly-efficient open-sourced library for GNNs. THU-DH (Gao et al., 2022) is an open-source PyTorch-based toolbox designed for general HNNs. It supports various hypergraph pre-processing methods (e.g., sampling, expansion) and convolution operators, facilitating the evaluation of HNNs.
Despite the progress made by these benchmarks, their inherent limitations render it impossible to conduct a comprehensive and fair comparison across various HNN algorithms. Firstly, Existing benchmarks primarily focus on node- and edge-level datasets, while overlooking graph-level datasets. Moreover, the node-level datasets they employ are predominantly homophilic. For example, HyFER conducts benchmark evaluation using only four homophilic academic networks. Secondly, the algorithmic coverage of current benchmarks is limited. For instance, HyFER implements only three HNN models, while THU-DH supports only a subset of methods proposed before 2022, thus excluding many recent SOTA approaches. Thirdly, Existing benchmarks mainly focus on model effectiveness (e.g., accuracy), while overlooking other critical aspects such as efficiency and complexity. Additionally, robustness and fairness are essential for trustworthy HNN applications, yet are rarely evaluated. Finally, existing studies adopt inconsistent experimental protocols, including variations in hyperparameter settings, initialization procedures, and convergence criteria, which can affect reproducibility and the reliability of cross-study comparisons. To address these issues, we introduce DHG-Bench, a more comprehensive and fair benchmark for HNNs, aiming to advance the state of the art and provide deeper insights into the challenges and opportunities of deep hypergraph learning.
D.2 Limitations.
DHG-Bench has several limitations that we aim to address in our furture work.
Our current benchmark is limited to a specific set of tasks, which might not reflect the full potential and versatility of HNN models. By incorporate novel benchmark tasks (e.g., hyperedge disambiguation (Kim et al., 2023), edge-dependent node classification (Choe et al., 2023), and hyperlink prediction (Tang et al., 2025)), we can gain a deeper understanding of the progress in the field and provide insights into how different algorithms perform across diverse applications.
Due to resource constraints and the availability of implementations, we could not include some of the latest state-of-the-art HNNs in our benchmark. This might impact the comprehensiveness of our evaluation, as some promising methods are not represented. We aim to address this by continuously updating our package and incorporating these algorithms as they become available.
Our current benchmark primarily focuses on the semi-supervised learning paradigm. However, task-specific labels are often scarce in real-world scenarios, and deep models are prone to overfitting when trained with such limited supervision signals (Li et al., 2024a). In future evaluations, we plan to incorporate recently proposed hypergraph self-supervised learning methods (Wei et al., 2022; Lee & Shin, 2023; Li et al., 2024a; Kim et al., 2024a; Lee et al., 2024) to provide insights into HNNs’ robustness in label-scarce settings and their ability to learn transferable representations without heavy reliance on supervision.

### v2

@@ A.2 Benchmark Algorithms
Figure A1 illustrates 17 HNN algorithms integrated into our DHG-Bench, including 10 spectral-based, 5 spatial-based, and 2 tensor-based methods. We introduce these methods in detail below.
A.2.1 Spectral-Based Algorithms
@@
HyperGCN (Yadati et al., 2019): HyperGCN approximates each hyperedge of the hypergraph by a set of pairwise edges connecting the vertices of the hyperedge, and treats the learning problem as a graph learning task on the approximated graph.
@@
HyperND (Prokopchik et al., 2022): HyperND develops a nonlinear diffusion process on hypergraphs that propagates both features and labels along the hypergraph structure. The novel diffusion incorporates a broad class of nonlinearities to increase the modeling capability, and the limiting point serves as a node embedding from which we make predictions with a linear model.
@@
SheafHyperGNN (Duta et al., 2023): SheafHyperGNN introduces a cellular sheaf framework for hypergraphs, enabling the modeling of complex dynamics while preserving their higher-order connectivity. Then, it generalizes the two commonly used hypergraph Laplacians to incorporate the richer structure sheaves offer and constructs two powerful neural networks capable of inferring and processing hypergraph sheaf structure.
HJRL (Yan et al., 2024): HJRL introduces a novel cross expansion method, which transforms both hypervertices and edges of a hypergraph to vertices in a standard graph. Then, a joint learning model is proposed to embed both hypervertices and hyperedges into a shared representation space. In addition, the algorithm employs a hypergraph reconstruction objective to preserve structural information in the model.
DPHGNN (Saxena et al., 2024): DPHGNN is a hybrid framework designed for effective feature representation in resource-constrained hypergraph settings. It introduces equivariant operator learning to capture lower-order semantics by inducing topology-aware inductive biases. It employs a dual-layered feature update mechanism: a static update layer provides spectral biases and relational features, while a dynamic update layer fuses explicitly aggregated features from the underlying topology into the hypergraph message-passing process.
@@
A.2.2 Spatial-Based Algorithms
HNHN (Dong et al., 2020): HNHN is a hypergraph convolution network with nonlinear activation functions applied to both hypernodes and hyperedges, combined with a normalization scheme that can flexibly adjust the importance of high-cardinality hyperedges and high-degree vertices depending on the dataset.
@@ A.2.2 Spatial-Based Algorithms
AllSetTransformer (Chien et al., 2022): AllSetTransformer is a novel HNN paradigm that implements each layer as a composition of two multiset functions. By incorporating the Set Transformer (Lee et al., 2019) into its architecture, it achieves greater modeling flexibility and enhanced expressive power.
ED-HNN (Wang et al., 2023a): ED-HNN is an architecture designed to approximate any continuous, permutation-equivariant hypergraph diffusion operator. The model is efficiently implemented by combining the star expansion (bipartite representation) of hypergraphs with standard message-passing neural networks, and supports scalable training via shared weights across layers.
HyperGT (Liu et al., 2024): HyperGT is a Transformer-based HNN architecture designed to capture global correlations among nodes and hyperedges. To preserve local structural information, it incorporates incidence-matrix-based positional encoding and a structure regularization term. These designs enable comprehensive hypergraph representation learning by jointly modeling global interactions and local connectivity patterns.
A.2.3 Tensor-Based Algorithms
@@
T-HyperGNN (Wang et al., 2024): T-HyperGNN is a general framework that integrates tensor hypergraph signal processing (t-HGSP) (Pena-Pena et al., 2023) to encode hypergraph structures using tensors. It models node interactions through multiplicative interaction tensors, elevating aggregation from traditional linear operations to higher-order polynomial mappings, thereby enhancing expressive power. To ensure scalability, T-HyperGNN introduces tensor-message-passing by exploiting tensor sparsity, enabling efficient processing of large hypergraphs with computational and memory costs comparable to matrix-based HNNs.
In addition, we include MLP and two GNN-based methods, CEGCN and CEGAT (Chien et al., 2022), as baselines in our comparative study. Both CEGCN and CEGAT are expansion-based approaches that transform a hypergraph into a pairwise graph via clique expansion (Zhou et al., 2006), where each hyperedge is converted into a clique over its incident nodes. Specifically, CEGCN applies GCN (Kipf & Welling, 2017) to the expanded graph, while CEGAT employs GAT (Veličković et al., 2018) to model node importance within the cliques.
Appendix B Details of the Experimental Settings
@@ B.1 General Experimental Settings
| SheafHyperGNN | init_hedge | rand, avg |
| sheaf_pred_block | MLP_var1, MLP_var2, MLP_var3, cp_decomp | |
@@
| TF-HNN | mlp_hidden_size | 64, 128, 256, 512, 1024 |
| # layers of classifier | 1, 2, 3, 4 | |
@@
| ED-HNN | alpha | 0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9 |
| # layers of | 0, 1, 2, 3 | |
| # layers of | 0, 1, 2, 3 | |
| # layers of | 0, 1, 2, 3 | |
@@
| # layers of module | 1, 2, 3, 4 | |
| # layers of module | 1, 2 | |
| # layers of module | 1, 2 | |
| HyperGT | attention_heads | 1, 2, 4, 8, 16 |
@@
We strive to follow the original implementations of various HNN methods from their respective papers or source codes and integrate them into a unified training and evaluation framework. All parameters are randomly initialized. We use the cross-entropy loss function (Mao et al., 2023) for all three benchmark classification tasks. Adam optimizer (Kingma, 2014) is adopted with an appropriate learning rate and weight decay to achieve the best performance on the validation split. Detailed hyperparameter settings and experimental environments are provided in Appendix B.2 and Appendix B.3, respectively. For evaluation, we follow prior studies in choosing task-specific metrics: accuracy for node classification (Feng et al., 2019; Chien et al., 2022; Wang et al., 2023a); AUROC (area under the ROC curve) and AP (average precision) for hyperedge prediction (Hwang et al., 2022; Ko et al., 2025; Yu et al., 2025; Tang et al., 2025); and both accuracy and Macro-F1 score for hypergraph classification (Feng et al., 2024). Higher values of these metrics indicate better predictive performance. In addition, to assess algorithmic fairness, we adopt two commonly used group fairness metrics: demographic parity () (Dwork et al., 2012) and equalized odds () (Hardt et al., 2016), with detailed definitions provided in Appendix B.5. For each method and dataset, we record the mean results and the standard deviation across 5 runs.
B.2 Hyperparameter Setting
We carefully tune hyperparameters to ensure a rigorous and unbiased evaluation of the integrated HNN methods. For algorithms without explicit hyperparameter guidelines in their original papers or source code, we perform a grid search with a reasonable budget across all datasets to identify optimal configurations. The search spaces are provided in Table A4. For detailed interpretations, please refer to the corresponding papers, and the complete hyperparameter configurations are available in our publicly released GitHub repository.
B.3 Experimental Environment
@@
B.4 Robustness Evaluation Settings
In our robustness study, we simulate data perturbation scenarios from three perspectives: structure, feature, and supervision signal. Each perturbation setting is repeated 5 times with different random seeds to account for randomness, and we report the average results. Our experiments primarily focus on the node classification task. The detailed experimental setups are as follows.
Structure-level Robustness Evaluation Setting. To analyze structure-level robustness, following (Cai et al., 2022), we randomly remove or add a proportion of node–hyperedge connections (i.e., hyperlinks) in the original hypergraph and then train and evaluate HNN algorithms on the perturbed structures. The modification ratio ranges from 0 to 0.9 to simulate varying levels of noise intensity.
Feature-level Robustness Evaluation Setting. To study feature-level robustness, we simulate two realistic types of feature perturbations: feature noise and feature sparsity. For feature noise, following (Wu et al., 2020), we add independent Gaussian noise to each feature dimension of all nodes with gradually increasing amplitude. Specifically, we use the mean of the maximum feature value of each node as the reference amplitude , and add Gaussian noise to each feature dimension, where and denotes the feature noise ratio. We evaluate model performance as varies from 0 to 0.9 with a step size of 0.1. For feature sparsity, following (Li et al., 2023), we randomly mask a certain proportion of node features by filling them with zeros, with the sparsity ratio ranging from 0 to 0.9 at an interval of 0.1.
Supervision-level Robustness Evaluation Setting. We study supervision-level robustness by simulating realistic noise and sparsity scenarios. For label noise, following (Dai et al., 2021), a certain proportion of training samples are randomly assigned incorrect labels by uniformly flipping them to one of the other classes. The noise ratio varies from 0 to 0.2 in increments of 0.05. Sparsity is introduced by reducing the ratio of training nodes, with the sparsity rate ranging from 0 to 0.8 with a step size of 0.2.
B.5 Fairness Evaluation Metrics
For fairness evaluation, we adopt two widely used group fairness metrics: demographic parity (DP) (Dwork et al., 2012), and equalized odds (EO) (Hardt et al., 2016). We focus on a binary classification task, with target label and binary sensitive attribute .
Demographic Parity. If the predicted result is independent of sensitive attributes , i.e., , then we can consider demographic parity is achieved. Formally, this criterion can be expressed as:
If a model satisfies demographic parity, the acceptance rate of different protected groups is the same. The deviation measure in the quantitative evaluation is given by:
where a smaller value indicates a fairer prediction distribution across groups.
Equalized Odds. If the predicted outcome and the sensitive attribute are conditionally independent given the ground-truth label , i.e., , then we consider equalized odds is achieved. The formula for this criterion is as follows:
If a model achieves equalized odds, the True Positive Rate (TPR) and False Positive Rate (FPR) are equal across different protected groups. The deviation measure is calculated as:
where a smaller value reflects more equitable predictive behavior across sensitive groups under the same ground-truth condition.
Appendix C Supplementary Experimental Results
C.1 Experimental Results on Effectiveness Evaluation
Table A5 shows the node classification results of all HNN algorithms on three datasets: NTU2012, ModelNet, and Ratings.
Table A6, A7 reports the full result of hyperedge prediction and hypergraph classification, respectively. Tensor-based methods are not considered in the hypergraph classification task, as they lack flexibility in supporting multi-graph training.
| Method | NTU2012 | ModelNet40 | Ratings |
| MLP | 88.591.27 | 96.880.23 | 28.470.76 |
| CEGCN | 84.931.12 | 92.340.24 | 26.651.61 |
| CEGAT | 84.141.77 | 92.020.26 | 28.230.50 |
| HGNN | 90.130.89 | 97.430.20 | 28.050.28 |
| HyperGCN | 75.784.82 | 91.153.88 | 27.340.72 |
| HCHA | 90.531.00 | 97.680.16 | 28.330.34 |
| LEGCN | 89.820.91 | 96.820.24 | 28.210.50 |
| HyperND | 88.981.56 | 97.180.58 | 28.320.38 |
| PhenomNN | 88.780.67 | 98.280.18 | 28.490.41 |
| SheafHyperGNN | 90.810.58 | 98.300.19 | 28.350.57 |
| HJRL | 88.151.18 | 96.330.30 | 26.900.55 |
| DPHGNN | 84.771.06 | 97.190.17 | 28.571.07 |
| TF-HNN | 91.690.75 | 98.380.11 | 28.560.68 |
| HNHN | 87.271.53 | 97.300.27 | 27.290.70 |
| UniGNN | 89.860.44 | 98.420.08 | 28.390.64 |
| AllSetTransformer | 90.171.03 | 98.070.21 | 27.321.11 |
| ED-HNN | 91.450.70 | 98.510.15 | 28.380.31 |
| HyperGT | 86.002.05 | 96.830.17 | 26.580.33 |
| EHNN | 87.990.39 | 97.970.17 | 28.950.81 |
| T-HyperGNN | 89.151.09 | 97.760.34 | 24.631.22 |
| Method | Cora | PubMed | Cora-CA | DBLP-CA | Actor | Pokec | ||||||
| AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | AUROC | AP | |
| MLP | 68.011.23 | 71.321.13 | 66.000.44 | 69.210.61 | 71.151.73 | 72.801.27 | 69.190.19 | 70.660.36 | 54.752.29 | 53.631.56 | 69.692.56 | 69.073.03 |
| CEGCN | 66.102.43 | 65.502.85 | 60.143.97 | 60.253.60 | 67.273.78 | 71.232.69 | 64.061.11 | 65.071.69 | 50.020.01 | 50.050.02 | 73.032.76 | 70.082.95 |
| CEGAT | 72.480.52 | 71.020.64 | 62.206.25 | 61.635.99 | 69.811.13 | 70.381.33 | 66.508.80 | 65.298.21 | 56.345.33 | 56.365.07 | 81.010.43 | 79.611.60 |
| HyperGT | 60.684.46 | 63.024.00 | 64.380.58 | 67.790.59 | 65.992.48 | 69.662.20 | 74.270.24 | 72.900.17 | 65.181.60 | 63.240.53 | 81.375.38 | 82.735.83 |
| MLP | 91.701.02 | 91.431.09 | 95.731.86 | 95.721.84 | 63.621.69 | 56.983.93 | 75.120.70 | 71.100.74 | 52.340.55 | 51.600.68 | 57.251.81 | 52.884.57 |
| CEGCN | 91.501.55 | 90.481.42 | 98.630.73 | 98.650.77 | 62.661.82 | 55.313.58 | 75.060.76 | 68.981.67 | 48.163.87 | 47.033.79 | 54.665.66 | 42.162.71 |
| CEGAT | 88.701.71 | 88.431.72 | 98.800.61 | 98.830.59 | 63.511.54 | 56.974.83 | 74.122.69 | 68.614.73 | 49.514.71 | 46.854.93 | 57.322.59 | 38.222.54 |
| SheafHyperGNN | 96.001.38 | 95.961.32 | 99.730.33 | 99.740.30 | 62.342.06 | 56.473.49 | 77.001.14 | 72.781.17 | 53.112.39 | 52.562.74 | 56.492.51 | 51.434.42 |
| HJRL | 96.100.80 | 95.980.85 | 99.600.53 | 99.570.52 | 63.092.83 | 56.543.62 | 77.821.47 | 73.731.92 | 51.843.52 | 51.133.29 | 57.102.79 | 44.197.19 |
| HyperGT | 91.601.42 | 91.291.53 | 96.271.93 | 96.281.88 | 61.494.32 | 55.146.02 | 73.821.27 | 69.361.45 | 54.471.33 | 51.551.08 | 54.352.72 | 47.495.11 |
C.2 Experimental Results on Robustness Evaluation
C.3 Experimental Results on Fairness Evaluation
Table A8 presents the full experimental results of fairness evaluation in terms of three metrics: accuracy (Acc), demographic parity (), and equalized odds ().
| Method | German | Bail | Credit | ||||||
| Acc | Acc | Acc | |||||||
| CEGCN | 69.602.78 | 6.194.59 | 6.465.48 | OOM | OOM | OOM | OOM | OOM | OOM |
| CEGAT | 69.122.44 | 9.004.77 | 8.523.82 | OOM | OOM | OOM | OOM | OOM | OOM |
| HJRL | 69.923.46 | 3.522.70 | 3.051.82 | OOM | OOM | OOM | OOM | OOM | OOM |
| ED-HNN | 70.163.15 | 4.063.05 | 4.072.75 | 94.260.77 | 8.050.64 | 1.510.26 | OOM | OOM | OOM |
| HyperGT | 68.882.01 | 5.052.88 | 4.362.59 | 94.330.62 | 7.681.13 | 1.641.37 | 79.830.39 | 4.172.50 | 2.691.97 |
@@ Appendix D Further Discussions
D.1 Related Works
Hypergraph neural networks (HNNs) (Yadati et al., 2019; Prokopchik et al., 2022; Wang et al., 2023a; Xie et al., 2025) have been promising tools for handling learning tasks involving higher-order data, with notable applications in various fields, such as social network analysis (Sun et al., 2023), bioinformatics (Li et al., 2025a), and recommender systems (Li et al., 2025b). However, there exists no established benchmark specifically dedicated to comprehensively evaluating hypergraph neural networks. In this section, we introduce a broader range of related studies concerning the comparative evaluations of HNNs, providing sufficient context for our benchmark work.
@@ D.1 Related Works
Kim et al. (Kim et al., 2024) recently presented the first survey dedicated to HNNs, with an in-depth and step-by-step guide. The survey comprehensively reviews existing HNN architectures, training strategies, and applications, establishing a foundational understanding crucial for advancing the field of HNNs. To further understand the expressive power of HNNs, Wang et al. (Wang et al., 2025) conduct the first theoretical analysis on the generalization performance of distinct HNN architectures, offering practical guidance for improving HNNs’ effectiveness. Nevertheless, systematic empirical evaluations of different HNN algorithms remain scarce, leaving a limited understanding of their comparative performance in practice. To facilitate the reproducibility and empirical evaluation of HNN algorithms, several open-sourced libraries have been developed in recent years. HyFER (Hwang et al., 2021) is a well-modularized framework for implementing and evaluating HNNs, dividing the entire learning process into data, model, and task components. Moreover, to address the scalability problem that most existing implementations suffer from, HyFER is built on top of Deep Graph Library (DGL) (Wang et al., 2019), which is a highly efficient open-sourced library for GNNs. DHG (Gao et al., 2022) is an open-sourced PyTorch-based toolbox designed for general HNNs. It supports various hypergraph preprocessing methods (e.g., sampling, expansion) and convolution operators, facilitating the evaluation of HNNs. TopoX (Hajij et al., 2024) is a suite of Python packages for machine learning on topological domains. These packages enhance and generalize functionalities found in mainstream hypergraph computations and learning tools, enabling them on topological domains. TopoBench (Telyatnikov et al., 2024) is a modular Python library that standardizes benchmarking and accelerates research in Topological Deep Learning (TDL). It supports training and comparing Topological Neural Networks (TNNs) across diverse domains, including graphs, simplicial complexes, cellular complexes, and hypergraphs. However, these libraries do not fully cover the latest HNN algorithms, datasets, and evaluation tasks, and they provide only limited empirical results without offering an in-depth and comprehensive analysis of existing HNN methods.
To fill the gap, we develop DHG-Bench, the first comprehensive benchmark tailored explicitly for HNNs. Distinguished by its broad coverage, DHG-Bench spans a wide range of algorithms, datasets, and evaluation tasks, thereby establishing a standardized and versatile framework for deep hypergraph learning research. Moreover, it provides comprehensive and systematic empirical evaluations that uncover the strengths and limitations of different algorithms. By offering such in-depth quantitative analyses, our benchmark fosters deeper insights into the challenges and opportunities of HNNs, thereby advancing the state-of-the-art in this emerging field.
D.2 Future Directions
Drawing upon our empirical analyses in the main text, we point out some promising future directions for the deep hypergraph learning community.
Developing adaptive HNN methods for diverse datasets and tasks. Our experiments in Section 4.1 reveal that existing HNN architectures show substantial performance disparities across datasets and tasks, limiting their applicability in diverse scenarios. Future research should focus on designing adaptive HNN architectures and training techniques that can better accommodate the unique characteristics of datasets from different domains and varying task granularities, thereby enhancing the generalization ability of HNNs.
Improving the efficiency of HNN methods. Observations in Section 4.2 indicate that many advanced HNN methods fail to balance efficiency and predictive performance, and often run out of memory on large-scale datasets. As the size of hypergraphs continues to grow exponentially, a key area of future research is the reduction of memory and computational complexity in HNN algorithms while maintaining satisfactory model utility. Inspired by the favorable efficiency–effectiveness trade-off achieved by TF-HNN, it would be promising to devise more powerful decoupled architectures specifically tailored for HNN.
Developing more robust HNN methods. Our experimental results in Section 4.3 show that HNN algorithms are affected by different types of data perturbations and are particularly vulnerable to those at the feature and supervision levels. Future work should emphasize enhancing the robustness of HNNs to resist varying degrees of data noise and even adversarial attacks, thereby ensuring reliable performance in a wide range of industrial applications.
Developing fairness-aware HNN methods. Empirical evidence in Section 4.4 suggests that HNNs are more prone to biased predictions than traditional MLPs. Future research should investigate the theoretical mechanisms through which high-order message passing exacerbates fairness issues and then develop fairness-aware HNN methods that mitigate such discriminatory behavior. Progress in this direction is essential to ensure the safe adoption of HNNs in high-stakes real-world applications such as crime prediction and credit evaluation.
Appendix E Package
We have developed DHG-Bench 333https://github.com/Coco-Hut/DHG-Bench, an open-sourced package that provides a comprehensive and unbiased platform for evaluating HNN algorithms and supporting future research in this domain.
As shown in Figure A5, the code structure is well-designed to ensure fair experimental setups across different algorithms, easy reproduction of the experimental results, and support for flexible assembly of models for experiments. The DHG-Bench consists of the following four key modules. The Config module includes the files that define the necessary hyperparameters and settings. The Data module is used to load and preprocess datasets. The Algorithm module has 17 built-in state-of-the-art algorithms, covering three representative categories: spectral-based, spatial-based, and tensor-based methods. The evaluation module supports multi-faceted testing of algorithmic performance, encompassing effectiveness, efficiency, robustness, and fairness.